It takes van twice as long as it takes Jun to do a certain piece of work. Working together they can do the work in 6 days. How long would it take Jun to do it alone?

Answer by nerdybill(7384)

It takes van twice as long as it takes Jun to do a certain piece of work. Working together they can do the work in 6 days. How long would it take Jun to do it alone?
.
Let J = time (days) it takes Jun to do it alone
then
2J = time (days) it takes Van
.
6(1/J + 1/(2J)) = 1
multiplying both sides by 2J:
6(2 + 1) = 2J
6(3) = 2J
18 = 2J
9 days = J (time it takes Jun)
.
Van:
2J = 2(9) = 18 days